# Smart Data Portfolios:

### v1

@@ 6.3 Comparing Portfolio Implications Across Use Cases
As the table shows, SDPs can provide the telecom sector with increased governance discipline analogous to the well established financial supervision frameworks. They can also improve explainability by bridging the gap between high-level regulatory principles and operational data selection. Although illustrated in telecom, the same governance logic extends to any domain where heterogeneous data categories carry differentiated risk, including finance, insurance, health, logistics, and mobility services.
@@ 7 Conclusion
Artificial intelligence systems increasingly mediate access to credit, connectivity, health, education, and public services. Yet prevailing governance approaches continue to focus on models and outputs, while the selection and justification of training data remain largely heuristic. The Smart Data Portfolio (SDP) framework addresses this structural gap by providing a model-agnostic, input-level governance and explainability layer, translating abstract regulatory requirements into measurable and auditable constraints on data allocation, using a portfolio logic familiar from financial supervision.
SDPs quantify input governance as an explicit information-risk trade-off. Each governed data category contributes jointly to informational return and to governance-adjusted risk, and varying portfolio weights traces an attainable region whose upper envelope defines the governance-efficient frontier. Regulators shape this region by specifying admissible categories, governance weight bands, and a policy risk cap, while institutions optimize within these boundaries using their preferred models and validation procedures. These choices are disclosed through Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports, which translate portfolio-level decisions into auditable explanations for regulators and intelligible explanations for affected individuals. By separating model design from input governance, SDPs enable stable, model-agnostic accountability aligned with long-standing supervisory practice in financial regulation.
The Smart Data Portfolio framework is designed for immediate use by actors who already bear responsibility for AI governance outcomes, including sector regulators, supervisory agencies, AI-governance consultancies, and in-house risk, compliance, and data-governance teams. It can be operationalized within current governance and MLOps workflows using existing data inventories, validation results, and documentation practices, without requiring new technical infrastructure or changes to model architectures. Even coarse category definitions and approximate governance scores are sufficient to identify dominated data allocations, make performance-risk trade-offs explicit, and prioritize remediation efforts. In this sense, SDPs function as a practical diagnostic and control instrument in the short term, while remaining fully compatible with future standardization and formal regulatory adoption.
The telecom illustration demonstrates how SDPs accommodate heterogeneous applications such as device finance, personalized recommendations, and network-quality prediction within a single governance architecture. Each use case is assigned its own governed data mixture while preserving a common optimization logic and reporting grammar. Across applications, SDPs allow regulators to evaluate input choices without inspecting proprietary model architectures, clarifying how governance risks accumulate and how institutions justify decisions through input-level explanations rather than internal model mechanics.
The policy implications are immediate. Article 10 of the EU AI Act on data governance already requires that training, validation, and testing data for high-risk systems satisfy rigorously documented quality and representativeness criteria (EUAIAct2024). SDP concepts can be incorporated into this regime through sectoral data taxonomies, reference risk-weight schedules, and governance weight bands defined in implementing acts, harmonized standards, or codes of practice (OECD2019; NIST2023). In the United States, the NIST AI Risk Management Framework highlights data and input management as a core dimension of trustworthy AI, and sector regulators such as the FCC, Ofcom, and BEREC are beginning to scrutinize AI-mediated communications and automated decision systems in telecom and adjacent markets (FCC2023; EDPB2018). Requiring Data Portfolio Statements and Data Portfolio Cards alongside existing model documentation would establish a common governance vocabulary that links high-level principles to enforceable input-selection rules.
Smart Data Portfolios provide a pathway to coherent public accountability by reframing explainability as an input-level property, allowing institutions and regulators to justify automated decisions through governed data allocations rather than opaque internal model mechanics. Like any portfolio model, SDPs rely on metric choice and category delineation. As a result, future work should therefore focus on pilot implementations, empirical estimation of governance-efficient frontiers, and robustness analysis under realistic institutional constraints. Overall, SDPs show that input governance can be structured as a quantitative, optimization-based discipline that scales across sectors, complements model-level tools, and supports transparent and accountable AI deployment.
@@
This appendix outlines several supervisory mechanisms that can be integrated into Smart Data Portfolios without altering the underlying optimization structure. Each mechanism has a direct analogue in long-standing financial-risk governance and provides regulators with additional levers for shaping data allocation. They are intentionally illustrative rather than exhaustive and are not required for baseline SDP implementation.
@@ A.1 Concentration limits for connected data sources
where $\theta_{k}$ is regulator-set. These linear constraints complement the weight-band structure in the main text and guard against systemic dependence on high-risk intermediaries (BaselLE2023).
@@ A.2 Scenario-based governance stress testing
A stress-based admissibility condition requires $\sigma^{(s)}_{p}(w)\leq\bar{\sigma}^{(s)}$ for all $s\in\mathcal{S}$. This approach exposes latent fragility in ways not captured by average-case risk measures (DFAST2024).
@@ A.3 Risk-weighted data exposures (D-RWA)
Supervisors may summarise portfolio exposure using a single linear index analogous to bank risk-weighted assets (RWA):
@@ A.4 Governance baselines via risk-parity weights
A natural baseline for benchmarking institutional portfolios is an equal-risk-contribution allocation (Maillard2010), defined as the solution to

### v2

@@ 6.3 Comparing Portfolio Implications Across Use Cases
As the table shows, SDPs provide a structured governance discipline that links regulatory objectives to operational data selection decisions. They can also improve explainability by bridging the gap between high-level regulatory principles and operational data selection. Although illustrated in telecom, the same governance logic extends to any domain where heterogeneous data categories carry differentiated risk, including finance, insurance, health, logistics, and mobility services.
@@ 7 Conclusion
Artificial intelligence systems increasingly operate within regulatory environments that place explicit responsibility on institutions to govern the training data underlying automated decision-making. While contemporary frameworks require documentation, representativeness, and risk management of datasets, they provide limited operational guidance for translating these obligations into implementable data-selection practices. The Smart Data Portfolio framework addresses this institutional gap by introducing a model-agnostic approach to governing AI training data through structured allocation and measurable governance risk.
By conceptualizing data categories as risk-bearing informational assets, SDPs make input governance observable, comparable, and auditable. Portfolio-level quantities linking informational performance to governance exposure allow regulators to constrain aggregate risk while preserving institutional flexibility in model design and deployment. In this framework, explainability shifts from post hoc inspection of model internals toward transparent justification of governed data use. Reporting instruments such as Data Portfolio Statements, Data Portfolio Cards, and Consumer Portfolio Reports translate portfolio choices into accountability mechanisms intelligible to regulators, auditors, and affected individuals.
The framework is designed for integration within existing compliance and risk-management environments. Because SDPs operate at the level of data allocation rather than model architecture, they remain compatible with heterogeneous AI systems and evolving technical practices. Sectoral illustrations demonstrate how diverse applications can operate under shared governance principles while maintaining application-specific data portfolios.
More broadly, Smart Data Portfolios suggest that input governance can be treated as a supervisory discipline analogous to risk management in finance. As implementation of the EU Artificial Intelligence Act and related governance frameworks progresses, operational methods for managing training data will become increasingly necessary (European Parliament and Council of the European Union, 2024; National Institute of Standards and Technology, 2023). Future research should therefore examine empirical estimation of governance-efficient frontiers, institutional adoption pathways, and auditing procedures under real-world regulatory constraints. By providing a quantitative yet institutionally interpretable framework for governing training data, SDPs contribute toward scalable and transparent accountability in large-scale AI deployment.
@@
This appendix outlines several supervisory mechanisms that can be integrated into Smart Data Portfolios without altering the underlying optimization structure. Each mechanism has a conceptual analogue in long-standing financial-risk governance and provides regulators with additional levers for shaping data allocation. They are intentionally illustrative rather than exhaustive and are not required for baseline SDP implementation.
@@ A.1 Concentration limits for connected data sources
where $\theta_{k}$ is regulator-set. These linear constraints complement the weight-band structure in the main text and guard against systemic dependence on high-risk intermediaries (Basel Committee on Banking Supervision, 2014).
@@ A.2 Scenario-based governance stress testing
A stress-based admissibility condition requires $\sigma^{(s)}_{p}(w)\leq\bar{\sigma}^{(s)}$ for all $s\in\mathcal{S}$. This approach exposes latent fragility in ways not captured by average-case risk measures (Board of Governors of the Federal Reserve System, 2024).
@@ A.3 Risk-weighted data exposures (D-RWA)
Supervisors may summarize portfolio exposure using a single linear index analogous to bank risk-weighted assets (RWA):
@@ A.4 Governance baselines via risk-parity weights
A natural baseline for benchmarking institutional portfolios is an equal-risk-contribution allocation (Maillard et al., 2010), defined as the solution to